OpenStudy (no_ellephant):

lim x->infinity x^4 +2x^3-3x^2+x-2

OpenStudy (anonymous):

Not Factorable.

OpenStudy (turingtest):

as x gets very large the x^4 "overtakes" the other terms therefor the limit is positive infinity, or DNE

OpenStudy (zarkon):

I like to write it like this \[\lim_{x\to\infty}(x^4 +2x^3-3x^2+x-2)\] \[\lim_{x\to\infty}x^4\left(1 +\frac{2}{x}-\frac{3}{x^2}+\frac{1}{x^3}-\frac{2}{x^4}\right)\] butchering notation... \[=\infty(1+0-0+0-0)=\infty\cdot1=\infty\]

OpenStudy (turingtest):

is \[\infty\cdot 1\]even defined technically? I thought it was iindeterminate.

OpenStudy (anonymous):

^no it's not indeterminate

OpenStudy (turingtest):

...no you're right, I just realized